# A bottleneck model with shared autonomous vehicles: Scale economies and price regulations

### v1

@@ 2.3 Equilibrium conditions
In this study, we define an equilibrium as a state in which no commuter can reduce their commuting cost by unilaterally changing either their departure time or their travel mode. Consistent with the existing literature (e.g. Wu and Huang, 2014; van den Berg and Verhoef, 2016), the closed-form solution of the equilibrium is obtained by solving the two equilibrium conditions sequentially. First, for any given mode-choice pattern $(N_{n},N_{a})$, we characterize the departure-time choice equilibrium. This equilibrium determines commuters’ departure-time decisions and yields equilibrium commuting costs of each mode given that pattern. Using these mode-specific commuting costs, we solve the mode-choice condition under which each commuter chooses the travel mode that minimizes their commuting cost. [^4] This hierarchical structure reflects a two-stage decision process: commuters choose a travel mode first and then select a departure time. The procedure above solves these decisions backwards.
@@ 2.3.1 Departure-time choice condition
$\int_{t\in\mathbb{R}}n_{n}(t)=N_{n},\quad\int_{t\in\mathbb{R}}n_{a}(t)=N_{a},$ (12)
where $n_{n}(t)$ and $n_{a}(t)$ denote the number of NV commuters and SAV commuters who arrive at the destination at time $t$, respectively. $c_{n}^{*}(N_{n},N_{a})$, $c_{a}^{*}(N_{n},N_{a})$ represent the mode-specific commuting costs in the departure-time choice equilibrium when the mode-choice pattern is $(N_{n},N_{a})$.
Conditions (9) and (10) state the departure-time choice conditions for NV and SAV commuters, respectively. Condition (11) is the queueing delay condition at the bottleneck: when a queue is present, the total departure-flow rate at the bottleneck equals the bottleneck capacity $\mu$; otherwise, the departure-flow rate does not exceed capacity. Owing to the capacity effect of SAVs, their effective departure-flow rate is scaled by $\kappa$. Condition (12) is the flow-conservation condition for each mode.
@@ 2.3.3 Stability condition
Because the traffic state that satisfies the mode-choice conditions is not necessarily unique, we examine the stability of equilibria if multiple equilibria arise. Previous studies have shown that economies of scale can lead to the existence of multiple equilibria (Tabuchi, 1993; Pandey and Lehe, 2024). In these circumstances, it is crucial to identify which equilibrium is likely to be realized by analyzing how the traffic state evolves as commuters adjust their choices (Beckmann et al., 1956).
@@
(NS) $V(N_{a})=0\ \text{implies that $(N_{n},N_{a})$ is the equilibrium}.$ (17)
The PC property requires that, whenever the mode-choice pattern is out of rest, the covariance between the growth rate of the number of commuters choosing each mode and its payoff (the negative of its commuting cost) be positive. The NS property requires that every rest point of the evolutionary dynamics be precisely an equilibrium point. Specific examples include the best response (Gilboa and Matsui, 1991), the Brown-von Neumann-Nash (Brown and Neumann, 1951), and the Smith dynamics (Smith, 1984).
@@ 3.1 Departure time choice equilibrium
This is because the VOT of SAV commuters is scaled by $\theta<1$: the ratio of their marginal schedule cost to VOT, $\beta/\theta\alpha$, exceeds that for NV commuters, $\beta/\alpha$. Consistent with classical bottleneck models, commuters with a higher effective schedule cost choose arrival times that are closer to their desired arrival time.
@@
$c_{a}^{*}(N_{n},N_{a})=\cfrac{\beta\gamma}{\mu(\beta+\gamma)}(\theta N_{n}+\kappa N_{a})+\theta t_{f}+p.$ (19)
Eqs. (18) and (19) show that the VOT reduction parameter $\theta$ affects only the commuting cost of SAV commuters, whereas the capacity-expansion parameter $\kappa$ influences the commuting costs of both modes, including NVs. This difference arises because the capacity effect shortens the SAV rush-hour period, which in turn shifts destination-arrival times of NV commuters. Because of the temporal sorting property, the shorter the SAV rush-hour window, the closer NV commuters can arrive to their desired arrival time. Consequently, when the capacity effect strengthens (i.e., either because $\kappa$ decreases or the number of SAV commuters increases), NV schedule costs decrease and, thus, their commuting costs.
By contrast, a decrease in $\theta$ reduces the VOT of SAV commuters only; it reduces their free-flow travel time and queuing-delay costs but leaves NV commuters unaffected. These asymmetric impacts of $\kappa$ and $\theta$ play an important role in the social cost analyses in Section 4 and Section 5.
@@ 3.2 Mode choice equilibrium
We then derive the mode-choice equilibrium using the mode-specific commuting costs (18) and (19) under the departure-time choice equilibrium. In the remainder of this section, we present closed-form equilibria for each fare-setting scenario introduced in Section 2 (see Appendix A for the derivation details).
3.2.1 Marginal-cost pricing
We first focus on MC pricing by setting the fare equal to the marginal cost, $p=m$. Substituting this condition into Eqs. (18)–(19) and the equilibrium conditions (13)–(14), we derive the closed-form mode-choice equilibrium, as follows:
@@ 3.2.1 Marginal-cost pricing
$A:=\cfrac{\beta\gamma(1-\theta)}{(\beta+\gamma)\mu},\quad B:=\theta t_{f}+m-(t_{f}+F_{n}),$ (20)
@@
This proposition suggests that the share of SAVs increases in cities with heavy congestion and long travel times: the number of SAV commuters grows when (i) the population size $N$ is large; (ii) the capacity $\mu$ is small or the free-flow travel time $t_{f}$ is long; and (iii) the penalty parameters $\beta$ and $\gamma$ are large. This is because the VOT effect of SAVs increases the incentive to use them as travel times become longer.
Condition (21) rules out corner cases in which all commuters choose a single mode. $B$ denotes the difference in mode-specific fixed costs when the fare is set to MC. The inequality $B>0$ ensures that, even at this low fare, the fixed cost of an SAV exceeds that of an NV; hence, some commuters still choose NVs. Conversely, ANB>0AN-B>0 requires a population to be large enough for congestion delays: commuters have an incentive to choose SAVs. As shown later, these two inequalities are necessary for both modes to be used not only in the benchmark case but also under AC pricing and a natural monopoly. They are consistent with our aim of analyzing how price regulations affect the bimodal transportation system. We therefore impose the following assumption for the remainder of the paper:
@@
3.2.2 Average-cost pricing
We next consider AC pricing, under which the fare equals the average cost, $p=AC(N_{a})$. Combining this condition with the mode-choice equilibrium conditions, we derive the following proposition, which shows that AC pricing may give rise to multiple equilibria:
@@ 3.2.2 Average-cost pricing
$N_{a1}^{AC*}=\cfrac{AN-B-\sqrt{(AN-B)^{2}-4AF_{a}}}{2A},$ (25)
$N_{a2}^{AC*}=\cfrac{AN-B+\sqrt{(AN-B)^{2}-4AF_{a}}}{2A}.$ (26)
If Condition (23) does not hold, the equilibrium is unique, and the equilibrium number of SAV commuters is given by $N_{a0}^{AC*}=0$.
We see that SAVs are used only when Condition (23) holds. This condition does not hold, for example, when $N$ is small or $F_{a}$ is large. In such situations, the small population relative to the fixed cost increases the average operating cost, which in turn drives up fares, making SAVs less attractive to commuters.
@@
The stability result suggests that SAV adoption may be hindered if AC pricing is implemented at its early stages. Specifically, if AC pricing is implemented before the number of SAV commuters exceeds $N_{a1}^{AC*}$, the system converges to the stable zero-adoption equilibrium, $N_{a0}^{AC*}$. The mechanism lies in the positive feedback effect created by AC pricing between the number of SAV commuters and their commuting costs. Once the number of SAV commuters falls below the unstable equilibrium $N_{a1}^{AC*}$, the commuting cost for SAVs exceeds that for NVs, i.e., $c_{n}^{*}<c_{a}^{*}$. This is because under AC pricing, fewer commuters raise the regulated fares, thereby increasing the SAV commuting cost. Hence, given the evolutionary dynamics $V(N_{a})$ satisfying the PC property, commuters have an incentive to switch from SAVs to NVs and no incentive to switch back. As a result, the number of SAV commuters keeps falling, and the system ultimately converges to the stable equilibrium $N_{a0}^{AC*}$ in which no one uses SAVs.
Conversely, if AC pricing is implemented after the number of SAV commuters exceeds $N_{a1}^{AC*}$, the system converges to the high-adoption equilibrium $N_{a2}^{AC*}$. This occurs because the positive feedback effect now drives adoption upward: as the number of SAV commuters increases, the fare decreases, reducing the SAV commuting cost below that of NVs, i.e., $c_{n}^{*}>c_{a}^{*}$. NV commuters thus have an incentive to switch to SAVs under the evolutionary dynamics, which further increases the SAV share. Therefore, the timing of implementing AC pricing should be carefully coordinated with the current level of SAV adoption.
Figure 2: Mode-specific commuting costs and the number of SAV commuters; solid and dashed circles represent the stable and unstable equilibria, respectively. $N=1,000$, $\kappa=0.01$, $\theta=0.7$, $\beta=0.4$, $\gamma=0.4$, $\mu=0.025$, $t_{f}=10$, $F_{n}=1$, $m=100$.
@@ 3.2.3 Natural monopoly
We then derive the equilibrium number of SAV commuters $N^{m*}_{a}$ by combining the demand function and the optimal fare. Moreover, substituting $N^{m*}_{a}$ and $p^{m*}$ into (4), we can identify the condition under which the service provider can gain profit. These results are summarized in the following proposition:
@@
Suppose that Condition (23) holds. Then, the profit $\pi$ becomes non-negative at the equilibrium. The equilibrium number of SAV commuters is expressed as follows:
@@ 3.3.1 Mode choice and cost patterns
This section investigates the effects of the price regulations by comparing the equilibria. We first obtain the following theorem about the mode choice and commuting-cost patterns:
@@
Suppose that Condition (23) holds. Then, the equilibrium number of SAV commuters under each price regulation satisfies the following relationship:
@@
Theorem 1 shows that commuting costs decrease as the number of SAV commuters increases. It further shows that MC pricing maximizes the SAV share and is therefore most beneficial for commuters; however, it imposes financial deficits on the service provider. When Condition (23) holds, both AC pricing and the natural-monopoly fare can reduce commuting costs relative to the costs incurred when no SAVs are used, without imposing deficits on the provider. Moreover, the number of SAV commuters at the high-adoption equilibrium under AC pricing, $N_{a2}^{AC*}$, exceeds the number at the monopoly equilibrium, $N_{a}^{m*}$. Accordingly, if such a superior equilibrium is achieved, the price regulation yields a larger reduction in commuting costs while maintaining the provider’s financial balance.
Interestingly, the theorem suggests that if AC pricing is implemented appropriately, the SAV share can achieve the high-adoption equilibrium $N_{a2}^{AC*}$ from any arbitrary initial state. Based on the theorem and the stability analysis under AC pricing, we propose the following two-step strategy for the case in which Condition (23) holds strictly:
@@
No price regulation. Keep fares unregulated and allow a natural monopoly.
@@
Activate AC pricing. Once the traffic state reaches the monopolistic equilibrium, impose AC pricing.
This strategy overcomes the drawback of AC pricing—its inability to encourage SAV adoption at a low market share—by temporarily allowing a natural monopoly. As scale economies lead to a natural monopoly, the (monopolistic) service provider sets the fare at $p^{m*}$ and the number of SAV commuters reaches $N^{m*}_{a}$. As shown in Eq. (30), $N^{m*}_{a}$ exceeds $N_{a1}^{AC*}$, the unstable equilibrium under AC pricing. Thus, imposing AC pricing at this point shifts the system toward the high-adoption equilibrium $N_{a2}^{AC*}$, thereby further increasing the SAV share and reducing commuting cost. Consequently, such a pricing strategy tailored to the diffusion stage facilitates the widespread adoption of SAVs in an economically viable manner.
@@ 3.3.2 Effects of capacity expansion
We further analyze the effect of capacity expansion on commuting costs when both NVs and SAVs are used. The sensitivity coefficient of the equilibrium commuting cost $c^{*}$ with respect to $\mu$ can be obtained by differentiating Eq. (18) as follows:
$\cfrac{\partial c^{*}}{\partial\mu}=\cfrac{\beta\gamma}{(\beta+\gamma)\mu}\left[-\cfrac{N-(1-\kappa)N_{a}}{\mu}-(1-\kappa)\cfrac{\partial N_{a}}{\partial\mu}\right].$ (34)
The first term represents the direct effect of the capacity expansion on the commuting cost, given the current mode choice pattern. The second term captures the indirect effect, wherein capacity expansion alters the SAV share, which in turn affects the commuting cost.
Because the first term is always negative, the sign of this sensitivity is governed by the second term. Noting that $\partial A/\partial\mu<0$ and $\partial B/\partial\mu=0$, we obtain $\partial N_{a}/\partial\mu>0$ under MC pricing or the natural-monopoly equilibrium. The sensitivity $\partial c^{*}/\partial\mu$ is therefore negative in these situations. By contrast, under AC pricing, the sign of $\partial N_{a}/\partial\mu$ is not predetermined. If $\partial N_{a}/\partial\mu<0$, the second term becomes positive, which may cause $\partial c^{*}/\partial\mu$ to turn positive as well.
Substituting the equilibrium number of SAV commuters under each price regulation, we obtain the following theorem:
@@
Under MC pricing and natural monopoly, the equilibrium commuting costs always decrease with capacity expansion. Under AC pricing, suppose that Condition (23) holds with strict inequality and consider the stable high-adoption equilibrium $N_{a2}^{AC*}$. The corresponding equilibrium commuting cost increases when the following condition holds:
$N-(1-\kappa)\left(\cfrac{N}{2}+\cfrac{(AN-B)AN-2AF_{a}}{2A\sqrt{(AN-B)^{2}-4AF_{a}}}\right)<0$ (35)
@@
Figure 3: The Downs–Thomson paradox under AC pricing. $N=250$, $\kappa=0.01$, $\theta=0.7$, $\beta=0.4$, $\gamma=0.4$, $t_{f}=10$, $F_{n}=1$, $m=100$.
This theorem states that AC pricing may cause the Downs–Thomson paradox in SAV systems because as capacity expands, the SAV share can decrease, i.e., $\partial N_{a}/\partial\mu<0$. Although capacity expansion uniformly reduces queuing delays for all commuters, the resulting cost savings are greater for NV commuters than for SAV commuters due to the difference in the VOT. This cost difference incentivizes SAV commuters to switch to NVs, reducing the SAV share. This reduction is further amplified under AC pricing since the fare rises with a declining SAV share: that is, capacity expansion triggers the negative feedback loop. As a result, the positive impact of capacity expansion diminishes, leading to higher commuting costs for all commuters. Figure 3 confirms the occurrence of this paradox under AC pricing.
@@
These findings underscore the importance of integrating hard and soft transport policies to provide commuters with convenient and financially sustainable transport options. Expanding capacity is not inherently inappropriate for improving traffic conditions under a natural monopoly; however, it may be ultimately ineffective if AC pricing is later implemented. Infrastructure development should be planned with a long-term perspective that considers the potential direction of future transport policies.

### v2

@@ 2.3 Equilibrium conditions
We define an equilibrium as a state in which no commuter can reduce their commuting cost by unilaterally changing either their departure time or their travel mode. Consistent with the existing literature (e.g. Wu and Huang, 2014; van den Berg and Verhoef, 2016), the closed-form solution of the equilibrium is obtained by solving the two equilibrium conditions sequentially. First, for any given mode-choice pattern $(N_{n},N_{a})$, we characterize the departure-time choice equilibrium. This equilibrium determines commuters’ departure-time decisions and yields equilibrium commuting costs of each mode given that pattern. Using these mode-specific commuting costs, we solve the mode-choice condition under which each commuter chooses the travel mode that minimizes their commuting cost. [^6] This hierarchical structure reflects a two-stage decision process: commuters choose a travel mode first and then select a departure time. The procedure above solves these decisions backwards.
@@ 2.3.1 Departure-time choice condition
$\color[rgb]{0,0,0}{\int_{t\in\mathbb{R}}n_{n}(t)\mathrm{d}t=N_{n},\quad\int_{t\in\mathbb{R}}n_{a}(t)\mathrm{d}t=N_{a}},$ (12)
where $n_{n}(t)$ and $n_{a}(t)$ denote the arrival flow rates of NV and SAV commuters at the destination at time $t$, respectively. $c_{n}^{*}(N_{n},N_{a})$, $c_{a}^{*}(N_{n},N_{a})$ represent the mode-specific commuting costs in the departure-time choice equilibrium when the mode-choice pattern is $(N_{n},N_{a})$.
Conditions (9) and (10) state the departure-time choice conditions for NV and SAV commuters, respectively. Condition (11) is the queueing delay condition at the bottleneck: when a queue is present, the total departure-flow rate at the bottleneck equals the bottleneck capacity $\mu$; otherwise, the departure-flow rate does not exceed capacity. [^7] While we do not model the congestion impact of SAV deadheading (empty-vehicle travel to the next pickup) for analytical tractability, it may be captured in reduced form as an effective reduction in bottleneck capacity (i.e., a lower effective $\mu$) if its impact is approximately stable over the peak period. Owing to the capacity effect of SAVs, their effective departure-flow rate is scaled by $\kappa$. Condition (12) is the flow conservation condition for each mode.
@@ 2.3.3 Stability condition
Because the traffic state that satisfies the mode choice conditions is not necessarily unique, we examine the stability of equilibria if multiple equilibria arise. Previous studies have shown that scale economies can lead to the existence of multiple equilibria (Tabuchi, 1993; Pandey and Lehe, 2024). In these circumstances, it is crucial to identify which equilibrium is likely to be realized by analyzing how the traffic state evolves as commuters adjust their choices (Beckmann et al., 1956).
@@
(NS) $V(N_{a})=0\ \text{if and only if $(N_{n},N_{a})$ is the equilibrium}.$ (17)
The PC property requires that, whenever the mode choice pattern is out of rest, the covariance between the growth rate of the number of commuters choosing each mode and its payoff (the negative of its commuting cost) be positive. The NS property requires that every rest point of the evolutionary dynamics be precisely an equilibrium point. Specific examples include the best response (Gilboa and Matsui, 1991), the Brown-von Neumann-Nash (Brown and Neumann, 1951), and the Smith dynamics (Smith, 1984).
@@ 3.1 Departure time choice equilibrium
This is because the VOT for SAV commuters is scaled by $\theta<1$: the ratio of their marginal schedule cost to VOT, $\beta/(\theta\alpha)$, exceeds that for NV commuters, $\beta/\alpha$. Consistent with classical bottleneck models, commuters with a higher effective schedule cost choose arrival times that are closer to their desired arrival time.
@@
$c_{a}^{*}(N_{n},N_{a})=\cfrac{\beta\gamma}{\mu(\beta+\gamma)}(\theta N_{n}+\kappa N_{a})+\theta t_{f}+p+w.$ (19)
Eqs. (18) and (19) show that the VOT reduction parameter $\theta$ affects only the commuting cost for SAV commuters, whereas the capacity-expansion parameter $\kappa$ influences the commuting costs for both NV and SAV commuters. This difference arises because the capacity effect shortens the SAV rush-hour period, thereby shifting the destination arrival times of NV commuters. Because of the temporal sorting property, the shorter the SAV rush-hour window, the closer NV commuters can arrive at their desired arrival time. Consequently, when the capacity effect strengthens (i.e., either because $\kappa$ decreases or because the number of SAV commuters increases), NV schedule costs decrease, which reduces their commuting costs.
By contrast, a decrease in $\theta$ reduces the VOT for SAV commuters only; it reduces their free-flow travel time and queuing-delay costs but leaves NV commuters unaffected. These asymmetric impacts of $\kappa$ and $\theta$ play an important role in the social cost analyses in Section 4 and Section 5.
@@ 3.2 Mode choice equilibrium
We then derive the mode choice equilibrium using the mode-specific commuting costs (18) and (19) under the departure time choice equilibrium. In the remainder of this section, we present closed-form equilibria for each fare-setting scenario introduced in Section 2 (see Appendix A for the derivation details).
3.2.1 Marginal cost pricing
We first focus on MC pricing by setting the fare equal to the marginal cost, $p=m$. Substituting this condition into Eqs. (18)–(19) and the equilibrium conditions (13)–(14), we derive the closed-form mode choice equilibrium, as follows:
@@ 3.2.1 Marginal cost pricing
$A:=\cfrac{\beta\gamma(1-\theta)}{(\beta+\gamma)\mu},\quad B:=\theta t_{f}+m+w-(t_{f}+F_{n}),$ (20)
@@
This proposition suggests that the share of SAV commuters increases in cities with heavy congestion and long travel times: $N_{a}$ increases when (i) the population size $N$ is large; (ii) the capacity $\mu$ is small or the free-flow travel time $t_{f}$ is long; and (iii) the penalty parameters $\beta$ and $\gamma$ are large. This is because the VOT effect of SAVs increases the incentive to use them as travel times become longer.
Condition (21) rules out corner cases in which all commuters choose a single mode. $B$ denotes the difference in mode-specific fixed costs when the fare is set to MC. The inequality $B>0$ ensures that, even at this low fare, the fixed cost of an SAV exceeds that of an NV; hence, some commuters still choose NVs. The condition $AN>B$ requires a population to be large enough for congestion delays: commuters have an incentive to choose SAVs. As shown later, these two inequalities are necessary for both modes to be used not only in the benchmark case but also under AC pricing and unregulated monopoly pricing. They are consistent with our aim of analyzing how fare regulations affect the bimodal transportation system. We therefore impose the following assumption for the remainder of the paper:
@@
3.2.2 Average cost pricing
We next consider AC pricing, under which the fare equals the average cost, $p=AC(N_{a})$. Combining this condition with the mode choice equilibrium conditions, we derive the following proposition, which shows that AC pricing may give rise to multiple equilibria:
@@ 3.2.2 Average cost pricing
$N_{a1}^{AC*}=\cfrac{AN-B-K}{2A},$ (25)
$N_{a2}^{AC*}=\cfrac{AN-B+K}{2A},$ (26)
where $K=\sqrt{(AN-B)^{2}-4AF_{a}}$. If Condition (23) does not hold, the equilibrium is unique, and the equilibrium number of SAV commuters is given by $N_{a0}^{AC*}=0$.
We see that SAVs are used only when Condition (23) holds. This condition does not hold, for example, when $N$ is small or $F_{a}$ is large. In such situations, the small population relative to the fixed cost increases the average operating cost, which in turn drives up the fare, making SAVs less attractive to commuters.
@@
The stability result suggests that SAV adoption may be hindered if AC pricing is implemented at its early stages. Specifically, if AC pricing is implemented before the number of SAV commuters $N_{a}$ exceeds $N_{a1}^{AC*}$, the system converges to the stable zero-adoption equilibrium, $N_{a0}^{AC*}$. The mechanism lies in the positive feedback effect created by AC pricing between the number of SAV commuters and their commuting costs. Once the number of SAV commuters falls below the unstable equilibrium $N_{a1}^{AC*}$, the commuting cost for SAV commuters exceeds that for NV commuters, i.e., $c_{n}^{*}<c_{a}^{*}$. This is because under AC pricing, fewer commuters raise the regulated fare, thereby increasing the SAV commuting cost. Hence, given evolutionary dynamics $V(N_{a})$ satisfying the PC property, commuters have an incentive to switch from SAVs to NVs and no incentive to switch back. As a result, the number of SAV commuters keeps falling, and the system ultimately converges to the stable equilibrium $N_{a0}^{AC*}$ in which no one uses SAVs.
Conversely, if AC pricing is implemented after $N_{a}$ exceeds $N_{a1}^{AC*}$, the system converges to the high-adoption equilibrium $N_{a2}^{AC*}$. This occurs because the positive feedback effect drives $N_{a}$ upward: as $N_{a}$ increases, the fare decreases, reducing the commuting cost for SAV commuters below that for NV commuters, i.e., $c_{n}^{*}>c_{a}^{*}$. NV commuters thus have an incentive to switch to SAVs under the evolutionary dynamics, which further increases $N_{a}$. Therefore, the timing of implementing AC pricing should be carefully coordinated with the current level of SAV adoption.
Figure 2: Mode-specific commuting costs and the number of SAV commuters; solid and dashed circles represent the stable and unstable equilibria, respectively. $N=1000$, $\kappa=0.01$, $\theta=0.7$, $\beta=0.4$, $\gamma=0.4$, $\mu=0.2$, $t_{f}=10$, $F_{n}=100$, $m=100$, $w=20$, $F_{a}=36000$.
@@ 3.2.3 Unregulated monopoly pricing
We then derive the equilibrium number of SAV commuters $N^{m*}_{a}$ by combining the demand function with the optimal fare. Substituting $N^{m*}_{a}$ and $p^{m*}$ into (4) yields the condition under which the service provider makes a profit. These results are summarized in the following proposition:
@@
Suppose that Condition (23) holds. Then, the profit $\pi$ in (4) becomes non-negative at the equilibrium. The equilibrium number of SAV commuters is expressed as follows:
@@ 3.3.1 Mode choice and cost patterns
This section compares equilibrium outcomes under three fare-setting scenarios. We first establish the following theorem on the ordering of the number of SAV commuters and associated commuting costs across the three scenarios:
@@
Suppose that Condition (23) holds. Then, the equilibrium number of SAV commuters under each fare-setting scenario satisfies the following relationship:
@@
Theorem 1 shows that commuting costs decrease as the number of SAV commuters increases. It further shows that MC pricing maximizes the SAV share and is therefore most beneficial for commuters; however, it imposes financial deficits on the service provider. When Condition (23) holds, both AC pricing and unregulated monopoly pricing can reduce commuting costs compared with those in the no-SAV case, while keeping the provider’s profit non-negative. In particular, the number of SAV commuters at the high-adoption equilibrium under AC pricing, $N_{a2}^{AC*}$, exceeds the number at the unregulated monopoly equilibrium, $N_{a}^{m*}$. Therefore, when the high-adoption equilibrium is reached, AC fare regulation achieves a larger reduction in commuting costs while maintaining the provider’s financial balance.
Interestingly, the theorem suggests that, with an appropriate timing of AC pricing, the system can converge to the high-adoption equilibrium $N_{a2}^{AC*}$ from any initial state. Based on the theorem and the stability analysis under AC pricing, we propose the following two-step pricing strategy for the case in which Condition (23) holds strictly:
@@
No fare regulation. Keep the fare unregulated and allow monopoly pricing.
@@
Introduce AC fare regulation. Once the traffic state reaches the monopoly equilibrium, implement AC pricing.
This strategy overcomes a drawback of AC pricing—its inability to promote SAV adoption from a low adoption level—by temporarily allowing unregulated monopoly pricing. As scale economies lead to a natural monopoly, the service provider sets the fare at $p^{m*}$, and the number of SAV commuters converges to $N^{m*}_{a}$. As shown in Eq. (30), $N^{m*}_{a}$ exceeds $N_{a1}^{AC*}$, the unstable equilibrium under AC pricing. Implementing AC pricing at this point therefore shifts the system toward the high-adoption equilibrium $N_{a2}^{AC*}$, further increasing the SAV adoption and reducing commuting cost. Consequently, this pricing strategy, contingent on the diffusion stage, facilitates widespread SAV adoption in an economically viable manner.
@@ 3.3.2 Effects of capacity expansion
We next examine how an increase in bottleneck capacity $\mu$ affects equilibrium commuting cost when both NVs and SAVs are used. The sensitivity of the equilibrium cost $c^{*}$ with respect to $\mu$ can be obtained by differentiating Eq. (18) as follows:
$\cfrac{\mathrm{d}c^{*}}{\mathrm{d}\mu}=\cfrac{\beta\gamma}{(\beta+\gamma)\mu}\left[-\cfrac{N-(1-\kappa)N_{a}}{\mu}-(1-\kappa)\cfrac{\mathrm{d}N_{a}}{\mathrm{d}\mu}\right].$ (34)
The first term represents the direct effect of capacity expansion and is always negative: for a given mode share, expanding capacity reduces congestion and the equilibrium commuting cost. The second term captures an indirect effect: the change in the equilibrium commuting cost shifts commuters’ mode choices, i.e., the equilibrium number of SAV commuters. This changes the aggregate capacity effect induced by SAVs, thereby further affecting the equilibrium commuting cost.
Under MC pricing and unregulated monopoly pricing, capacity expansion always reduces the equilibrium commuting cost, even though it induces a mode shift from SAVs to NVs. Increasing capacity reduces congestion-related delays for both modes; however, because NV commuters have a higher VOT than SAV commuters, the resulting cost reduction is larger for NV commuters. This induces some commuters to switch from SAVs to NVs, implying $\mathrm{d}N_{a}/\mathrm{d}\mu<0$ and a positive indirect-effect term in Eq. (34). Nevertheless, this indirect effect never outweighs the direct effect in both fare-setting scenarios, and thus the total derivative remains negative. This can be confirmed by substituting the closed-form solutions in Eqs. (22) and (29) into Eq. (34):
[MC pricing] $\cfrac{\mathrm{d}N^{MC*}_{a}}{\mathrm{d}\mu}=-\cfrac{B}{A\mu}\quad$ $\Rightarrow\quad$ $\cfrac{\mathrm{d}c^{MC*}}{\mathrm{d}\mu}=-\cfrac{\beta\gamma}{\beta+\gamma}\cfrac{\kappa N}{\mu^{2}}<0.$ (35)
[Monopoly pricing] $\cfrac{\mathrm{d}N^{m*}_{a}}{\mathrm{d}\mu}=-\cfrac{B}{2A\mu}\quad$ $\Rightarrow\quad$ $\cfrac{\mathrm{d}c^{m*}}{\mathrm{d}\mu}=-\cfrac{\beta\gamma}{\beta+\gamma}\cfrac{(1+\kappa)N}{2\mu^{2}}<0.$ (36)
By contrast, under AC pricing, the derivative in Eq. (34) can be positive: the Downs–Thomson paradox, where capacity expansion increases commuting costs, can occur, particularly in the high-adoption equilibrium. Substituting the equilibrium number of SAV commuters under AC pricing into Eq. (34) yields the following theorem.
@@
Under AC pricing, suppose that Condition (23) holds with strict inequality and consider the stable high-adoption equilibrium $N_{a2}^{AC*}$. The equilibrium commuting cost increases as capacity increases when the following condition holds:
$K<\left(AN-B-\cfrac{2F_{a}}{N}\right)\cfrac{1-\kappa}{1+\kappa}.$ (37)
@@
Figure 3: Downs–Thomson paradox under AC pricing. $N=250$, $\kappa=0.01$, $\theta=0.7$, $\beta=0.4$, $\gamma=0.4$, $t_{f}=10$, $F_{n}=500$, $m=499$, $w=100$, $F_{a}=10500$.
The paradox can arise because, in the high-adoption equilibrium under AC pricing, capacity expansion can trigger a positive feedback loop that amplifies the shift from SAVs to NVs. As under MC and monopoly pricing, capacity expansion makes NVs relatively more attractive, reducing $N_{a}$ and making the indirect term positive. Under AC pricing, this reduction in $N_{a}$ raises the fare $p=m+F_{a}/N_{a}$, which further discourages SAV use and reinforces the shift to NVs. As $N_{a}$ decreases, the aggregate capacity effect induced by SAVs weakens. Then, the indirect effect can outweigh the direct effect of capacity expansion, increasing the equilibrium commuting cost. Figure 3 confirms the occurrence of this paradox.
Eq. (37) indicates that the paradox is more likely to arise when (i) $\kappa$ is small and (ii) $K$ is small. When $\kappa$ is small, SAV commuters have a strong capacity advantage. In that case, a shift from SAV to NV can substantially worsen congestion because a decline in the SAV share causes a large reduction in the aggregate capacity effect induced by SAVs.
Second, when $K$ is small, the high-adoption equilibrium number of SAV commuters, $N_{a2}^{AC*}$, becomes highly sensitive to the bottleneck capacity $\mu$. This can be seen from the derivative with respect to $\mu$:
$\cfrac{\mathrm{d}N_{a2}^{AC*}}{\mathrm{d}\mu}=-\cfrac{N_{a2}^{AC*}(N-N_{a2}^{AC*})}{\mu(N_{a2}^{AC*}-N_{a1}^{AC*})}=-\cfrac{AN_{a2}^{AC*}(N-N_{a2}^{AC*})}{\mu K}$ (38)
A smaller $K$ therefore tends to magnify $\left|\mathrm{d}N_{a2}^{AC*}/\mathrm{d}\mu\right|$, strengthening the indirect effect.
To clarify this mechanism, we examine how much the equilibrium number of SAV commuters must adjust to restore $\Delta c\equiv c_{n}^{*}-c_{a}^{*}$ to zero after a capacity change creates a cost difference between the two modes. Differentiating $\Delta c\equiv c_{n}^{*}-c_{a}^{*}$ with respect to $N_{a}$ yields:
$\cfrac{\mathrm{d}\Delta c}{\mathrm{d}N_{a}}=\cfrac{F_{a}}{(N_{a})^{2}}-A.$ (39)
This derivative decomposes the effect of $N_{a}$ on the mode-cost difference $\Delta c$ into two terms. The first term $F_{a}/(N_{a})^{2}$ captures the fare effect under AC pricing: as $N_{a}$ increases, the fare $p=m+F_{a}/N_{a}$ decreases, which reduces $c_{a}^{*}$ and thereby increases $\Delta c$. The second term A-A captures the congestion and schedule-delay effect: as $N_{a}$ increases (and $N_{n}$ decreases), NV commuters face smaller congestion and schedule-delay costs, which decreases $\Delta c$. Importantly, the net effect of these two terms around the equilibrium is governed by $K$. Evaluating the derivative at the high-adoption equilibrium yields:
$\left.\cfrac{\mathrm{d}\Delta c}{\mathrm{d}N_{a}}\right|_{N_{a}=N_{a2}^{AC*}}=-\cfrac{K}{N_{a2}^{AC*}}$ (40)
Hence, when $K$ is close to zero, the two effects nearly offset each other and $\Delta c$ becomes locally insensitive to changes in $N_{a}$. As a result, a capacity-induced change in relative costs requires a large adjustment in $N_{a}$ to restore $\Delta c=0$, implying that $\left|\mathrm{d}N_{a2}^{AC*}/\mathrm{d}\mu\right|$ can be large. This strengthens the indirect effect and makes the paradox more likely.
@@
These findings underscore the importance of integrating hard and soft transport policies to provide commuters with convenient and financially sustainable transport options. Expanding capacity is not inherently inappropriate for improving traffic conditions under unregulated monopoly pricing; however, it may be ultimately ineffective if AC pricing is later implemented. Infrastructure development should be planned with a long-term perspective that considers the potential direction of future transport policies.